### v1

# If It’s Not Enough, Make It So: Reducing Authentic Data Demand in Face Recognition through Synthetic Faces

Abstract
Recent advances in deep face recognition have spurred a growing demand for large, diverse, and manually annotated face datasets. Acquiring authentic, high-quality data for face recognition has proven to be a challenge, primarily due to privacy concerns. Large face datasets are primarily sourced from web-based images, lacking explicit user consent. In this paper, we examine whether and how synthetic face data can be used to train effective face recognition models with reduced reliance on authentic images, thereby mitigating data collection concerns. First, we explored the performance gap among recent state-of-the-art face recognition models, trained with synthetic data only and authentic (scarce) data only. Then, we deepened our analysis by training a state-of-the-art backbone with various combinations of synthetic and authentic data, gaining insights into optimizing the limited use of the latter for verification accuracy. Finally, we assessed the effectiveness of data augmentation approaches on synthetic and authentic data, with the same goal in mind. Our results highlighted the effectiveness of FR trained on combined datasets, particularly when combined with appropriate augmentation techniques.
I Introduction
Face Recognition (FR) stands as one of the predominant biometric technologies in various applications, ranging from logical access control [26, 39, 4] to portable devices [9]. Existing FR models have reached impressive accuracy thanks to the recent advances in deep learning, including margin-penalty-based softmax losses [13, 39, 4, 26] and neural architectures [20, 8]. This development, however, has been made possible by the public availability of large-sized, identity-labeled, FR training databases [11, 19, 45].
Figure 1: Average verification accuracy on the testing dataset (y-axis) vs. number of authentic identities from CASIA-WebFace (WF) in the training dataset (x-axis). In all settings, the verification accuracy improved by increasing the number of authentic identities in the training dataset. Also, combining synthetic (black line, DCFace; green line, GANControl) with a limited subset of authentic datasets improved the FR performance, in comparison to the case where only a limited subset of authentic identities (red line) is used to train FR. The number of synthetic identities in the combined dataset experiments is fixed (10K).
Face databases include millions of images often gathered from the Internet without user consent, thereby giving rise to concerns on their legal and ethical use for FR development. Notably, the approval of the General Data Protection Regulation (GDPR) [1] by the European Union (EU) in 2018 has significantly amplified criticism concerning the privacy-related issues associated with the use of publicly available face datasets obtained without adequate consent. The GDPR explicitly affirms individuals’ ”right to be forgotten” and imposes more stringent regulations concerning the acquisition, distribution, and utilization of biometric data, making adherence to such regulations a key, intricate endeavor when applied to face datasets. Consequently, several databases frequently employed for FR training, including but not limited to [11, 3, 19], have been retracted by their creators, to prevent potential legal issues. These circumstances have raised doubts upon the continuity of FR research, as the availability of such fundamental resources has become uncertain.
Synthetic data has recently emerged as a promising alternative to genuine datasets for FR training [10], serving as a viable solution to address the prevailing legal and ethical concerns [34, 6, 8, 2, 28]. This trajectory of research has been significantly propelled by the advancement of Deep Generative Models (DGMs), which facilitate the generation of synthetic samples thanks to their ability to capture the probability distribution underlying a face dataset. This generative process can be constrained to various factors, including but not limited to age, facial expression, head pose, and lighting [14, 37, 38, 17]. The majority of DGM techniques employed for the creation of synthetic faces build upon Generative Adversarial Networks (GANs) [18, 14, 36, 37, 6] or Diffusion Models (DMs) [21, 33, 16].
Deep face recognition methods, such as [29, 8, 34], have increasingly leveraged GANs [14, 23] to generate synthetic data for FR training. However, state-of-the-art (SOTA) FR models trained on synthetic data have often exhibited a drop in verification accuracy, in comparison to FR models trained on authentic data. This discrepancy in performance primarily stems from the training datasets’ limited identity discrimination [6] or their minimal intra-class variance [8, 34]. DMs have emerged as a viable alternative to GANs for image synthesis, but this progress comes at the cost of stability and a substantial reduction in training performance. Unfortunately, several questions are still open on the extent to which authentic and synthetic data can be effectively combined to tackle the limitations inherent to both domains.
In this paper, our primary objective is to investigate the suitability of combined (authentic and synthetic) training datasets for model development in FR. This exploration can provide a better understanding of how to address both the performance inadequacies associated with synthetic datasets and the privacy-related concerns that characterize authentic datasets, thereby facilitating the creation of highly accurate FR models, with reduced authentic data demand. As teased in Figure 1, training FR models with a combination of synthetic data and subsets of authentic identities is visibly beneficial in terms of FR performance in comparison to cases where only limited subsets of authentic identities are available, as we assumed that the synthetic data could be generated without any limitation and that small numbers of authentic identities could be collected with proper user consent. Specifically, our novel contribution is threefold:
  • We investigated the impact of training models with different ratios of synthetic and authentic identities, under a fixed total number of training images.
  • We explored whether a training dataset with a fixed amount of synthetic identities and an increasing number of authentic ones could lead to (at least) comparable accuracy w.r.t. a large, authentic-only dataset.
  • We evaluated the influence of training data augmentation on accuracy, by augmenting (i) both the synthetic and authentic subsets and (ii) exclusively the synthetic subset, identifying and discussing the optimal settings.
II Related Work
With the ongoing success of deep generative models in generating high-quality and realistic face images [37, 15, 21, 33], several recent works proposed the use of synthetic data in FR development [34, 8, 2, 6, 5, 27, 10]. The synthetic data for FR training can be categorized based on the type of underlying generative model under three categories, GAN-Based [34, 8, 6, 7], digital rendering [2], and diffusion-based [30, 5, 27]
SynFace [34] and SFace [6] were among the earliest works that proposed the use of GAN-generated face images for FR training. To enhance intra-class diversity in synthetic data, SynFace proposed to generate synthetic data using an attribute-conditional GAN model, i.e., DiscoFaceGAN [15], and perform identity and domain mixup. SFace also analyzed the impact of Style-GAN [25] training under class conditional settings. Furthermore, the extent to which transferring knowledge from the pretrained model on authentic data improves the performance of synthetic-based FR was analyzed, going beyond FR training only on synthetic data. UsynthFace [8] leveraged unlabelled synthetic data for unsupervised FR training. The work also presented an extensive data augmentation method based on GANs [15] along with color and geometric transformation [12]. Conversely, ExFaceGAN [7] presented a framework to disentangle identity information in learned latent spaces of unconditional GANs, aiming at generating multiple samples of any synthetic identity.
DigiFace-1M [2] employed a digital rendering pipeline, leveraging facial geometry models, a diverse array of textures, hairstyles, and 3D accessories, along with robust data augmentation techniques during training. However, it comes at a considerable computational cost during the rendering process. DigiFace-1M also proposed to combine synthetic with authentic data during FR training to improve the verification accuracy of synthetic-based FR using a small and fixed number of authentic identities.
Very recently, IDiff-Face [5] and DCFace [27] adopted diffusion models to generate synthetic data for FR training, achieving SOTA verification accuracy for synthetic-based FR. Specifically, IDiff-Face included fizziness in the identity condition to induce variations in the generated data. DCFace proposed a two-stage generative framework. In the first stage, an image of a novel identity using an unconditional diffusion model is generated and an image style from the style bank is selected. In the second stage, the generated images and style are mixed using a dual conditional diffusion model. GANDiffFace [30] proposed to combine GAN and diffusion models to generate synthetic data for FR training that exhibit certain attributes such controllable demographic distribution.
Although these approaches [27, 5, 2, 6, 34] achieved relatively high verification accuracy when they were evaluated on authentic datasets, their performance is lower than the one obtained by FR models trained on authentic data [13]. To overcome this challenge, SFace [6] applied knowledge distillation on the embedding level between the model trained on the authentic data and the one trained on synthetic data. DigiFace [2] and DCFace [27] combined their synthetic datasets with a subset of the authentic CASIA-WebFace dataset [41], aiming at improving the verification accuracy of synthetic-based FR. Both methods demonstrated that synthetic-based verification accuracy can be improved using a limited set of authentic data. Wang et at. [40] proposed to combine synthetic data from SynFace [34] with a small and limited subset of authentic data. The work proposed to utilize the same backbone to extract features from synthetic and authentic identity and then learn multi-class classification through two classification layers for synthetic and authentic classes. However, these works [2, 27, 40] only utilized a subset of authentic data, with a fixed number of authentic identities, and they did not investigate the performance gain when different ratios of synthetic and authentic identities are combined in the FR training. Recently, Melzi et al. [31] organized a competition at WACV 2024, aiming to explore the use of synthetic data in FR training and to attract the development of solutions for synthetic-based FR. The challenge considered two main tasks, training FR only with synthetic data and training FR with both synthetic and authentic data.
Unlike previous research that combined fixed, (in terms of numbers of images and identities) sets of synthetic and authentic data [2, 27, 40, 31], we investigate in detail FR performances using different combinations of authentic and synthetic data. We first fixed the total number of identities in the training dataset and investigated the performance of different ratios of synthetic and authentic identities. Then, we investigated the FR performance using a fixed number of synthetic identities with different subsets of authentic data. Finally, we investigated the impact of data augmentation on synthetic and authentic subsets.
III Methodology
This work aims, at first, to investigate the performance gaps between FR models trained on synthetic datasets and the ones trained on privacy-sensitive authentic datasets. Then, we explored the performance gains of synthetic-based FR by including a limited set of authentic identities in the training dataset. Finally, we investigate the impact of data augmentation on synthetic-based FR verification accuracy.
III-A Datasets Preparation
The presented investigations in this paper are performed using four training datasets, two authentic and two synthetic.
As the authentic data, we adopted the well-known CASIA-WebFace [41] and MS1MV2 [13] datasets. CASIA-WebFace consists of 0.5M images of 10K identities. MS1MV2 is the refined version of the MS-Celeb-1M dataset [19]. MS1MV2 contains 5.8M images of 85K identities. To provide comparable results across FR trained on authentic-only data and on synthetic and authentic combinations, we utilized a subset of 10K identities of MS1MV2 in the conducted experiments. This subset will be noted as M2-S and contains the first 10K identities from MS1MV2.
The synthetic datasets come from GAN-Control [37] and DCFace [27]. Each of these datasets contains 0.5M images of 10K identities, i.e., 50 images per identity. The GAN-Control dataset is generated using an attribute conditional GAN-based model [37] trained on the FFHQ dataset [24]. Each synthetic identity in the GAN-Control dataset is generated by fixing the identity condition and randomizing the attribute conditions, i.e., age, pose, illumination and expression. From now on, the GAN-Control dataset will be noted as GC.
The second synthetic dataset is generated via DCFace [27]. The latter is based on two-stage diffusion models. In the first stage, a high-quality face image of a novel identity is generated using unconditional diffusion models [21] trained on FFHQ [24], with an image style randomly selected from a style bank. In the second stage, the generated images and styles from stage 1 are mixed using a dual conditional diffusion model [21] trained on CASIA-WebFace [41] to generate an image with a specific identity and style. The dataset generated by DCFace will be noted as DC. All images are aligned and cropped to $112\times 112$ using landmarks obtained by the Multi-task Cascaded Convolutional Networks (MTCNN) [42], following [13]. All the training and testing images are normalized to have values between -1 and 1.
Figure 2: FR Training Paradigm Overview. Subsets of authentic and synthetic data are combined to form the training dataset.
III-B Data Mixing
Given a synthetic dataset $S$ with a number of identities equal to $N$ and an authentic dataset $A$ with a number of identities equal to $M$, we combined these two datasets into one dataset $C=s\bigcup a$, where $s\subseteq S$ and $a\subseteq A$ are subsets of $S$ and $A$, respectively, with a total number of identities equal to $H$, as follows. In the first evaluation setting, we set $H$ to be equal to $M$ and sampled $m<M$ from the authentic dataset and $M-m<N$ from the synthetic dataset. We evaluated several $m\in[1\mathrm{K},2\mathrm{K},3\mathrm{K},4\mathrm{K},5\mathrm{K},6\mathrm{K},7\mathrm{K},8\mathrm{K},9\mathrm{K}]$. In this stage, we aim to boost the verification accuracy of synthetic-based FR by combining synthetic with authentic face images into one dataset. We fixed the total number of identities, in this case, to provide a fair comparison across the considered settings. In the second stage, we considered the whole synthetic dataset $S$ and combined it with a subset $a\subseteq A$. The sampled $a$ contains a number of identities equal to $1\mathrm{K}$, $2\mathrm{K}$, $4\mathrm{K}$, $6\mathrm{K}$, $8\mathrm{K}$ or $10\mathrm{K}$. In this scenario, we assumed that the synthetic data could be generated without any limitation and that a small number of authentic identities could be collected with proper user consent.
III-C Data Augmentation
Considering the constrained range of visual diversity in synthetic face data, in comparison to large and realistic appearance variations in authentic data [10, 8, 27], we evaluated in this work the impact of augmenting the FR training dataset on FR verification accuracy, as follows.
Baseline (no augmentation)
We refer to procedures commonly used for training FR models, which involve augmenting the training dataset solely through the application of a random horizontal flip, with a probability of $0.5$.
Random augmentation
We evaluated the impact of enriching the FR training dataset on the FR verification accuracy by employing random augmentation techniques [12, 8], denoted as RandAug. The augmentation space of RandAug includes the following color and geometric transformations [8]: horizontal-flipping, rotation, x-axis translation, y-axis translation, x-axis shear, y-axis shear, sharpness adjusting, auto-contrasting, adjusting contrast, solarizing, posterizing, equalizing, adjusting color, adjusting brightness, performing a resized crop, and converting to grayscale. RandAug includes two hyper-parameters, Q and M, to select the number of operations Q and the magnitude M of each transformation, respectively. We set M to 16 and Q to 4, following  [8]. In our analysis, we first experimented with the application of a random augmentation on both subsets of the training dataset, i.e., authentic and synthetic. In a later stage, we only augmented the synthetic subset (this approach is visually depicted in Figure 2).
III-D Model Training
We relied on FR models having the well-known ResNet-50 architecture [20] as the backbone and CosFace [39] as the loss function; the latter is defined as:
$L_{CosFace}=\frac{1}{N}\sum\limits_{i\in N}-log\frac{e^{s(cos(\theta_{y_{i}})-m)}}{e^{s(cos(\theta_{y_{i}})-m)}+\sum\limits_{j=1,j\neq y_{i}}^{c}e^{s(cos(\theta_{j}))}}$
where $c$ is the number of classes (identities), N is the batch size, $m$ is the margin penalty applied on the cosine angle $cos(\theta_{yi})$ between the feature representation $x_{i}$ of the sample $i$ and its class center $y_{i}$, $s$ is scale parameters. In all conducted experiments, the margin $m$ is set to 0.35 and the scale parameter $s$ to 64, following [39].
During the training, Stochastic Gradient Descent (SGD) is employed as an optimizer with an initial learning rate of 0.1. The learning rate is divided by 10 at 22, 30 and 40 training epochs [6]. The models have been trained for 40 epochs with a total batch size of 512.
III-E Evaluation Benchmarks
We evaluated the trained FR models on the well-known benchmark verification pairs that accompany the following datasets: Labeled Faces in the Wild (LFW) [22], AgeDB-30 [32], Celebrities in Frontal to Profile in the Wild (CFP-FP) [35], Cross-Age LFW [44], and Cross-Pose LFW [43]. Results for all benchmarks are reported as verification accuracy in %, thus adhering to their official evaluation protocol.
IV Experimental Results
Our experiments analyzed, at first, the performance gaps between FR models trained on authentic-only and synthetic-only data created with SOTA generative methods (Section IV-A). Then, we explored how combining different ratios of authentic and synthetic data can influence FR performance (Section IV-B) and whether the combination of (scarce) authentic and synthetic data can actually overcome the limitations of both data domains (Section IV-C). Finally, we explored the effects of the random augmentation technique, when applied to all training images and only to the synthetic ones (Section IV-D). Our experiments were conducted independently on the GC and DC synthetic datasets, without any data interchange and ensuring a clear separation between them. For each synthetic dataset, we incorporated varying percentages ($10\%,20\%,40\%,60\%,80\%,$ and $100\%$) of the authentic datasets’ identities.
IV-A Gap in FR training on Synthetic or Authentic data (RQ1)
In our initial analysis, we aimed to understand the performance disparities between FR models trained exclusively on authentic data and those exclusively trained on synthetic data generated through SOTA generative methods. This investigation sought to shed light on the impact of data source variation on model performance, while considering an equal number of identities across datasets.
To address this objective, Table I presents the verification accuracy for models trained exclusively on synthetic datasets (first rows group) and models trained with an increasing amount of authentic data from the CASIA-WebFace dataset only (second rows group). Consistent with previous research [2, 27], the model trained on the full authentic dataset showed the highest performance, with an average accuracy of $94.63\%$ across the five considered benchmarks. Notably, the model trained on DC data only (second row, average accuracy of $89.56\%$) surpassed the performance of the model trained on GC data only (first row, average accuracy of $79.38\%$).
IV-B Mixing Ratio Impact on Size-fixed Training Data (RQ2)
Train Data Id/Img. LFW ()($\uparrow$)( ↑ ) AgeDB-30 ()($\uparrow$)( ↑ ) CFP-FP ()($\uparrow$)( ↑ ) CA-LFW ()($\uparrow$)( ↑ ) CP-LFW ()($\uparrow$)( ↑ ) Avg ()($\uparrow$)( ↑ )
GC (10K) 10K / 50 93.22 77.60 73.03 82.25 70.80 79.38
DC (10K) 10K / 50 98.55 89.70 85.33 91.60 82.62 89.56
WF (1K) 1K / 45 95.60 81.97 80.81 84.78 76.83 83.99
WF (2K) 2K / 45 97.23 86.30 87.88 89.36 82.07 88.57
WF (3K) 3K / 45 98.62 90.32 90.04 90.83 84.30 90.82
WF (4K) 4K / 45 97.76 91.32 91.70 91.72 85.98 91.69
WF (5K) 5K / 45 99.15 92.75 93.24 92.40 87.38 92.98
WF (6K) 6K / 46 99.25 93.80 93.36 92.85 88.77 93.61
WF (7K) 7K / 46 99.37 93.58 94.07 93.13 89.05 93.84
WF (8K) 8K / 46 99.35 94.71 95.01 93.18 89.25 93.90
WF (9K) 9K / 46 99.42 94.32 95.03 93.73 89.70 94.44
WF (10K) 10K / 46 99.55 94.55 95.31 93.78 89.95 94.63
GC (1K) $\cup$ WF (9K) 10K / 46 99.43 94.62 95.10 93.12 90.02 94.46
GC (2K) $\cup$ WF (8K) 10K / 47 99.35 94.07 95.19 93.05 89.78 94.29
GC (3K) $\cup$ WF (7K) 10K / 47 99.35 93.58 94.90 92.98 89.73 94.11
GC (4K) $\cup$ WF (6K) 10K / 48 99.38 93.68 94.10 92.67 89.43 93.85
GC (5K) $\cup$ WF (5K) 10K / 48 99.30 93.42 93.34 92.60 89.07 93.55
GC (6K) $\cup$ WF (4K) 10K / 48 99.17 92.18 92.87 92.45 88.05 92.94
GC (7K) $\cup$ WF (3K) 10K / 49 98.93 91.93 91.97 91.82 86.98 92.33
GC (8K) $\cup$ WF (2K) 10K / 49 98.62 90.18 89.60 90.45 85.68 90.91
GC (9K) $\cup$ WF (1K) 10K / 50 98.17 89.15 87.49 88.78 83.30 89.38
DC (1K) $\cup$ WF (9K) 10K / 46 99.50 94.38 95.43 93.63 90.10 94.61
DC (2K) $\cup$ WF (8K) 10K / 47 99.45 94.65 95.56 93.55 90.18 94.68
DC (3K) $\cup$ WF (7K) 10K / 47 99.33 94.75 95.14 93.52 90.13 94,57
DC (4K) $\cup$ WF (6K) 10K / 48 99.43 94.28 95.01 93.42 90.45 94.52
DC (5K) $\cup$ WF (5K) 10K / 48 99.37 94.25 94.90 93.33 89.57 94.28
DC (6K) $\cup$ WF (4K) 10K / 48 99.27 94.10 94.41 93.33 89.52 94.13
DC (7K) $\cup$ WF (3K) 10K / 49 99.35 93.55 94.27 92.87 89.03 93.81
DC (8K) $\cup$ WF (2K) 10K / 49 99.17 93.27 93.77 92.68 88.10 93.40
DC (9K) $\cup$ WF (1K) 10K / 50 99.03 93.20 92.54 92.23 87.32 92.86
TABLE I: Verification accuracy on size-fixed training sets (authentic dataset: WF). First rows group: accuracy on full synthetic datasets; Second rows group: accuracy on different subsets of authentic WF; Third and fourth rows group: accuracy on different subsets of authentic WF and synthetic GC and DC datasets, respectively. It can be noticed that the accuracy of synthetic-based models improved by adding limited subsets of authentic data. Highest accuracy scores for each rows group in bold.
In our second analysis, considering a fixed total number of identities in the training dataset, we were interested in assessing whether using different combinations of authentic and synthetic data for training the considered models influences the resulting verification performance. Specifically, we set the total number of identities to $10\mathrm{K}$ and evaluated the FR performance using different ratios of synthetic and authentic identities in the training datasets.
The results presented in Table I (third and fourth rows groups) demonstrate that incorporating a limited number of authentic identities (e.g., $1\mathrm{K}$) resulted in notable improvements compared to models trained exclusively on synthetic identities only (first rows group). More precisely, by combining only $1\mathrm{K}$ authentic identities with $9\mathrm{K}$ identities from the DC dataset, we observed a performance improvement w.r.t. the respective fully synthetic baselines of $0.49\%$ on LFW, $3.90\%$ on AgeDB-30, $8.45\%$ on CFP-FP, $0.69\%$ on CA-LFW and $5.69\%$ on CP-LFW, with an overall average improvement of $3.68\%$ across datasets. Mixing the same $1\mathrm{K}$ authentic identities with the $9\mathrm{K}$ identities generated by the other method, GC, led to more noticeable improvements. We observed an improvement over the synthetic baselines of $5.31\%$ on LFW, $14.88\%$ on AgeDB-30, $20.35\%$ on CFP-FP, $7.94\%$ on CA-LFW and $17.66\%$ on CP-LFW, with an overall $12.60\%$ average improvement.
Given the performance differences between synthetic and authentic baselines, a consistent improvement was observed on all benchmarks as the number of involved authentic identities increase. Mixing GC data with the authentic data (third rows group) steadily improves the average performances from $89.38\%$ ($9\mathrm{K}$ synthetic and $1\mathrm{K}$ authentic identities) to $94.46\%$ ($1\mathrm{K}$ synthetic and $9\mathrm{K}$ authentic identities). This trend persisted while combining DC data with the authentic data (fourth rows group), with an increase in the average accuracy from $92.86\%$ to $94.68\%$.
Models trained on only subsets of authentic datasets, such as using $5\mathrm{K}$ authentic identities and $5\mathrm{K}$ synthetic ones, demonstrated highly competitive results (average accuracy of $99.30\%$ for GC and $94.28\%$ for DC) compared to models trained on the entire authentic dataset (average accuracy of $94.63\%$). Notably, certain combinations that involved DC identities even achieved superior performances than the full authentic-only baseline. Specifically, these observations were made for the settings with $2\mathrm{K}$ to $3\mathrm{K}$ synthetic identities for AgeDB-30, $1\mathrm{K}$ to $2\mathrm{K}$ synthetic identities for CFP-FP, and $1\mathrm{K}$ to $4\mathrm{K}$ synthetic identities for CP-LFW.
IV-C Performance Analysis as Authentic Data Increases (RQ3)
In our third analysis, we explored whether a training dataset with a large, fixed number of synthetic identities and a progressively growing number of authentic ones could achieve accuracy levels, at the very least, comparable to a larger dataset consisting solely of authentic identities, and we investigated the minimal ratio of authentic data required. To this end, Table II and III present the FR verification accuracy achieved by models trained on a large, fixed number of synthetic identities ($10\mathrm{K}$) and increasing amounts of authentic identities ($1\mathrm{K}$, $2\mathrm{K}$, $4\mathrm{K}$, $6\mathrm{K}$, $8\mathrm{K}$, $10\mathrm{K}$) from the WF and MS-2 datasets, respectively.
In both tables, we can observe that the accuracy becomes consistently higher by increasing the number of authentic identities in the training dataset, starting from a fixed number of synthetic identities from either GC or DC ($10\mathrm{K}$). This observation is evident when we compare the average accuracy estimates across rows groups in each table. By combining WF subsets with $10\mathrm{K}$ GC identities, models achieved an average accuracy ranging from $89.94\%$ ($1\mathrm{K}$ authentic identities) to $94.73\%$ ($10\mathrm{K}$ authentic identities). The same trend was observable also while combining WF subsets with DC identities, which led to an accuracy ranging from $93.09\%$ ($1\mathrm{K}$ authentic identities) to $95.06\%$ ($10\mathrm{K}$ authentic identities). However, it is worth noting that the accuracy obtained by combining the full $10\mathrm{K}$-identity synthetic dataset with $10\mathrm{K}$ authentic identities ($95.06\%$) were slightly worse than the ones obtained by combining it with $8\mathrm{K}$ authentic identities ($95.08\%$) from WF. Mixing M2-S with both GC and DC identities led to the same type of observation, reporting an accuracy ranging from $83.66\%$ to $95.08\%$ and from $90.72\%$ to $92.75\%$, respectively. As also visually emerging from Figure 3, all the settings involving combined datasets achieved a higher accuracy than the one of their respective synthetic-only baseline (first two row), with an improvement often of about $30\%$, making further comparisons under another granularity unnecessary.
To gain deeper insights into the influence of a specific type of generated data (GC or DC), we narrowed our focus to comparing the verification accuracy between combined and authentic-only datasets, while maintaining an identical number of authentic identities. This comparison was conducted within the same rows group in both Table II and III.
Train Data Id/Img. LFW ()($\uparrow$)( ↑ ) AgeDB-30 ()($\uparrow$)( ↑ ) CFP-FP ()($\uparrow$)( ↑ ) CA-LFW ()($\uparrow$)( ↑ ) CP-LFW ()($\uparrow$)( ↑ ) Avg ()($\uparrow$)( ↑ )
GC (10K) 10K / 50 93.22 77.60 73.03 82.25 70.80 79.38
DC (10K) 10K / 50 98.55 89.70 85.33 91.60 82.62 89.56
WF (1K) 1K / 45 95.60 81.97 80.81 84.78 76.83 83.99
GC (10K) $\cup$ WF (1K) 11K / 49 97.95 88.80 87.24 89.68 83.57 89.44
DC (10K) $\cup$ WF (1K) 11K / 49 99.13 93.45 92.83 92.43 87.63 93.09
WF (2K) 2K / 45 97.23 86.30 87.88 89.36 82.07 88.57
GC (10K) $\cup$ WF (2K) 12K / 49 98.67 90.45 90.90 91.50 85.92 91.49
DC (10K) $\cup$ WF (2K) 12K / 49 99.15 93.03 92.61 92.52 87.78 93.02
WF (4K) 4K / 45 97.76 91.32 91.70 91.72 85.98 91.69
GC (10K) $\cup$ WF (4K) 14K / 49 99.27 92.57 93.06 92.32 88.45 93.13
DC (10K) $\cup$ WF (4K) 14K / 49 99.45 94.47 94.64 93.20 89.58 94.27
WF (6K) 6K / 46 99.25 93.80 93.36 92.85 88.77 93.61
GC (10K) $\cup$ WF (6K) 16K / 49 99.38 93.30 94.39 91.12 89.43 93.52
DC (10K) $\cup$ WF (6K) 16K / 49 99.42 94.35 94.59 93.73 89.97 94.41
WF (8K) 8K / 46 99.35 94.71 95.01 93.18 89.25 93.90
GC (10K) $\cup$ WF (8K) 18K / 48 99.46 94.23 95.23 93.37 90.47 94.55
DC (10K) $\cup$ WF (8K) 18K / 49 99.55 95.15 95.88 93.83 90.98 95.08
WF (10K) 10K / 46 99.55 94.55 95.31 93.78 89.95 94.63
GC (10K) $\cup$ WF (10K) 20K / 48 99.52 94.53 95.26 93.82 90.53 94.73
DC (10K) $\cup$ WF (10K) 20K / 49 99.50 94.92 95.71 94.08 90.98 95.06
TABLE II: Verification accuracy on increasingly larger number of authentic identities in training sets (authentic dataset: WF). First group: accuracy of models trained on synthetic GC and DC datasets, each with 10K identities. First row in each group: accuracy of models trained by using different subsets of authentic WF only. Second and third rows in each group: accuracy of models obtained by combining different subsets of authentic WF with 10K synthetic identities from GC and DC datasets, respectively. It can be noticed that the accuracy of synthetic-based models improved by adding a limited subset of authentic data. Moreover, combining synthetic with authentic datasets (e.g., GC (10K) $\cup$ WF (10K)) led to higher accuracy, in comparison to the case where FR is only trained on authentic WF dataset/subsets. For each rows group, the highest accuracy score is highlighted in bold.
Train Data Id/Img. LFW ()($\uparrow$)( ↑ ) AgeDB-30 ()($\uparrow$)( ↑ ) CFP-FP ()($\uparrow$)( ↑ ) CA-LFW ()($\uparrow$)( ↑ ) CP-LFW ()($\uparrow$)( ↑ ) Avg ()($\uparrow$)( ↑ )
GC (10K) 10K / 50 93.22 77.60 73.03 82.25 70.80 79.38
DC (10K) 10K / 50 98.55 89.70 85.33 91.60 82.62 89.56
M2-S (1K) 1K / 45 93.15 79.85 74.49 81.62 69.55 79.73
GC (10K) $\cup$ M2-S (1K) 11K / 49 96.47 85.02 76.67 85.80 74.33 83.66
DC (10K) $\cup$ M2-S (1K) 11K / 49 98.63 90.92 89.10 92.18 82.78 90.72
M2-S (2K) 2K / 45 96.43 83.20 78.49 84.45 73.85 83.28
GC (10K) $\cup$ M2-S (2K) 12K / 49 98.00 86.87 81.26 88.05 78.72 86.58
DC (10K) $\cup$ M2-S (2K) 12K / 49 98.85 91.60 89.39 92.62 83.15 91.12
M2-S (4K) 4K / 45 98.53 91.42 87.24 92.33 83.43 90.59
GC (10K) $\cup$ M2-S (4K) 14K / 49 99.08 93.03 89.09 93.35 86.05 92.12
DC (10K) $\cup$ M2-S (4K) 14K / 49 99.02 92.57 90.01 93.00 84.48 91.82
M2-S (6K) 6K / 46 99.30 94.20 90.56 93.87 87.23 93.03
GC (10K) $\cup$ M2-S (6K) 16K / 49 99.45 94.37 91.66 94.05 87.98 93.50
DC (10K) $\cup$ M2-S (6K) 16K / 49 99.07 93.07 90.77 93.28 84.95 92.23
M2-S (8K) 8K / 46 99.32 95.07 90.31 94.22 87.08 93.20
GC (10K) $\cup$ M2-S (8K) 18K / 48 99.52 95.35 93.61 94.75 88.97 94.44
DC (10K) $\cup$ M2-S (8K) 18K / 49 99.22 93.45 91.66 93.40 85.53 92.65
M2-S (10K) 10K / 46 99.55 95.97 92.94 94.87 89.67 94.60
GC (10K) $\cup$ M2-S (10K) 20K / 48 99.63 95.83 94.64 95.07 89.82 95.00
DC (10K) $\cup$ M2-S (10K) 20K / 49 99.17 93.50 91.44 93.90 85.73 92.75
TABLE III: Verification accuracy on increasingly larger number of authentic identities in training sets (authentic dataset: M2-S). First group: accuracy of models trained on synthetic GC and DC datasets, each with 10K identities. First row in each group: accuracy of models trained by using different subsets of authentic M2-S only. Second and third rows in each group: accuracy of models obtained by combining different subsets of authentic M2-S with 10K synthetic identities from GC and DC datasets, respectively. It can be noticed that the accuracy of synthetic-based models improved by adding a limited subset of authentic data. Moreover, combining synthetic with authentic datasets (e.g., GC (10K) $\cup$ M2-S (10K)) led to higher accuracy, in comparison to the case where FR is only trained on authentic M2-S dataset/subsets. For each rows group, the highest accuracy score is highlighted in bold.
Figure 3: Verification accuracy on testing data (y-axis) vs. number of authentic identities in training data (x-axis). Top row refers to authentic data sampled from WF, bottom row to authentic data sampled from M2-S. The verification accuracy improved by increasing the number of authentic identities in the training dataset. Also, combining synthetic with a limited subset of authentic datasets (black and green lines) improved the FR performance in comparison to the case where only a limited subset of authentic identities (red lines) is used to train FR. The number of synthetic identities in the combined dataset experiments is fixed (10K). Plots in the top (bottom) row visually resemble the results reported in Table II (Table III).
Our results show, for small numbers of authentic identities, the improvement gained from adding the large synthetic dataset is more noticeable; however, it becomes less apparent when the number of authentic identities is higher. Specifically, training the models on DC (GC) data combined with authentic identities subsets from WF led, on average, to $10.83\%$ ($6.49\%$), $5.02\%$ ($3.30\%$), $2.81\%$ ($1.57\%$), $0.85\%$ ($-0.10\%$), $1.26\%$ ($0.69\%$) and $0.45\%$ ($0.11\%$) improvements w.r.t. the baseline trained only on authentic data. Each comparison was made by taking into consideration an authentic training subset involving 1K, 2K, 4K, 6K, 8K and 10K authentic identities, respectively. It also emerges that, under a given amount of authentic data, the impact of DC data is more evident than that of GC data.
Following the same protocol, the training of FR models on the full DC (GC) dataset combined with authentic identity subsets from M2-S led to improvements in accuracy by $13.78\%$ ($4.93\%$), $9.41\%$ ($3.96\%$), $1.36\%$ ($1.69\%$), $-0.86\%$ ($0.51\%$), $-0.59\%$ ($1.33\%$), and $-1.96\%$ ($0.42\%$). These results, for small numbers of authentic identities, confirm that the inclusion of the full synthetic dataset results in a substantial increase of accuracy and that DC identities are more effective that those from GC. However, differently from what we observed for WF authentic data, the performance on DC identities combined with 6K or more authentic identities from M2-S not only achieved lower performances compared to combining with GC identities, but also adversely affected performance with respect to using the authentic data only. This observation sheds a light on the importance of devising future work about methods for creating synthetic data functional to improving model accuracy, in an end-to-end manner.
IV-D Augmentation Impact on Combined Training Data (RQ4)
In our final analysis, we investigated the impact of data augmentation, as specified in Section III-C, on verification accuracy concerning combined datasets. Our experiments involved applying data augmentation to two distinct subsets of the training dataset: exclusively on synthetic identities and on the entire combined dataset. We compared the results from both scenarios with the ones that emerged from the previous research question, which will constitute the baselines in this subsection. To address this, Table IV and V illustrate the FR verification accuracy achieved by models trained on a fixed number of synthetic identities ($10\mathrm{K}$, from GC and DC respectively) combined with increasing amounts of authentic identities ($1\mathrm{K}$, $2\mathrm{K}$, $4\mathrm{K}$, $6\mathrm{K}$, $8\mathrm{K}$, $10\mathrm{K}$) from WF. The augmentation was applied to both authentic and synthetic subsets or only to the latter.
Our results show that applying data augmentation on both authentic and synthetic subsets (RAug-All settings) led to overall lower average accuracy, in comparison to the cases in which no data augmentation was applied. On average, models trained after augmenting all the identities part of the training dataset, when WF constituted the authentic component, were impacted by $-0.01\%$ ($+0.54\%$), $+0.23\%$ ($-0.43\%$), $-0.11\%$ ($-0.43\%$), $+0.21\%$ ($+0.10\%$), $-0.18\%$ ($-0.92\%$) and $-0.14\%$ ($-0.22\%$) w.r.t. the baselines trained on not augmented data. Each comparison was made by taking into consideration combined and authentic training sets involving 1K, 2K, 4K, 6K, 8K and 10K authentic identities, respectively. A slight but consistent improvement w.r.t. the combined baselines is instead observable after augmenting only the synthetic training subset (RAug-Synt settings). In this scenario, FR models trained on authentic identities subsets from WF combined with augmented synthetic samples from DC (GC) led to $-0.01\%$ ($+0.31\%$), $+0.49\%$ ($-0.30\%$), $+0.10\%$ ($+0.03\%$), $+0.27\%$ ($+0.52\%$), $-0.06\%$ ($0\%$) and $-0.06\%$ ($0.17\%$) changes in average verification accuracy.
Train Data Method LFW ()($\uparrow$)( ↑ ) AgeDB-30 ()($\uparrow$)( ↑ ) CFP-FP ()($\uparrow$)( ↑ ) CA-LFW ()($\uparrow$)( ↑ ) CP-LFW ()($\uparrow$)( ↑ ) Avg ()($\uparrow$)( ↑ )
GC (10K) $\cup$ WF (1K) - 97.95 88.80 87.24 89.68 83.57 89.44
GC (10K) $\cup$ WF (1K) RAug-All 98.27 88.87 89.31 89.35 83.82 89.92
GC (10K) $\cup$ WF (1K) RAug-Synt 98.25 88.47 88.76 89.68 83.47 89.73
GC (10K) $\cup$ WF (2K) - 98.67 90.45 90.90 91.50 85.92 91.49
GC (10K) $\cup$ WF (2K) RAug-All 98.80 89.88 90.52 90.67 85.65 91.10
GC (10K) $\cup$ WF (2K) RAug-Synt 98.83 90.61 89.60 91.30 85.75 91.22
GC (10K) $\cup$ WF (4K) - 99.27 92.57 93.06 92.32 88.45 93.13
GC (10K) $\cup$ WF (4K) RAug-All 99.02 92.23 92.69 92.07 87.67 92.74
GC (10K) $\cup$ WF (4K) RAug-Synt 99.10 92.28 93.03 92.45 88.92 93.16
GC (10K) $\cup$ WF (6K) - 99.38 93.30 94.39 91.12 89.43 93.52
GC (10K) $\cup$ WF (6K) RAug-All 99.37 92.80 94.33 92.78 88.81 93.62
GC (10K) $\cup$ WF (6K) RAug-Synt 99.35 93.13 94.77 92.75 89.97 93.79
GC (10K) $\cup$ WF (8K) - 99.46 94.23 95.23 93.37 90.47 94.55
GC (10K) $\cup$ WF (8K) RAug-All 99.32 93.25 94.47 92.35 89.10 93.70
GC (10K) $\cup$ WF (8K) RAug-Synt 99.48 94.25 95.13 93.48 90.40 94.55
GC (10K) $\cup$ WF (10K) - 99.52 94.53 95.26 93.82 90.53 94.73
GC (10K) $\cup$ WF (10K) RAug-All 99.48 93.97 95.27 93.53 90.38 94.53
GC (10K) $\cup$ WF (10K) RAug-Synt 99.42 94.37 95.40 93.57 90.60 94.67
TABLE IV: Impact of augmenting the training dataset on FR verification accuracies (authentic dataset: WF; synthetic dataset: GC). For different subsets of the authentic dataset combined with the synthetic dataset, we reported the results under three settings: no augmentation, both synthetic and authentic samples are randomly augmented (RAug-All) and only synthetic samples are augmented (RAug-Synt). For each group of experimental settings, the best average accuracy is in bold. The authentic identities are sampled from WF and the synthetic ones are from GC.
Train Data Method LFW ()($\uparrow$)( ↑ ) AgeDB-30 ()($\uparrow$)( ↑ ) CFP-FP ()($\uparrow$)( ↑ ) CA-LFW ()($\uparrow$)( ↑ ) CP-LFW ()($\uparrow$)( ↑ ) Avg ()($\uparrow$)( ↑ )
DC (10K) $\cup$ WF (1K) - 99.13 93.45 92.83 92.43 87.63 93.09
DC (10K) $\cup$ WF (1K) RAug-All 99.15 93.23 92.81 92.40 87.82 93.08
DC (10K) $\cup$ WF (1K) RAug-Synt 99.13 93.43 92.80 92.38 87.70 93.09
DC (10K) $\cup$ WF (2K) - 99.15 93.03 92.61 92.52 87.78 93.02
DC (10K) $\cup$ WF (2K) RAug-All 99.18 93.02 93.57 92.57 87.83 93.23
DC (10K) $\cup$ WF (2K) RAug-Synt 99.25 93.25 93.48 92.86 88.46 93.46
DC (10K) $\cup$ WF (4K) - 99.45 94.47 94.64 93.20 89.58 94.27
DC (10K) $\cup$ WF (4K) RAug-All 99.33 94.32 94.86 92.93 89.37 94.16
DC (10K) $\cup$ WF (4K) RAug-Synt 99.37 94.33 94.60 93.67 89.83 94.36
DC (10K) $\cup$ WF (6K) - 99.42 94.35 94.59 93.73 89.97 94.41
DC (10K) $\cup$ WF (6K) RAug-All 99.40 94.42 95.40 93.53 90.28 94.61
DC (10K) $\cup$ WF (6K) RAug-Synt 99.50 94.32 95.06 93.95 90.48 94.66
DC (10K) $\cup$ WF (8K) - 99.55 95.15 95.88 93.83 90.98 95.08
DC (10K) $\cup$ WF (8K) RAug-All 99.55 95.02 95.69 93.78 90.53 94.91
DC (10K) $\cup$ WF (8K) RAug-Synt 99.62 94.77 95.73 94.02 90.98 95.02
DC (10K) $\cup$ WF (10K) - 99.50 94.92 95.71 94.08 90.98 95.04
DC (10K) $\cup$ WF (10K) RAug-All 99.55 94.72 95.71 94.07 90.48 94.91
DC (10K) $\cup$ WF (10K) RAug-Synt 99.55 94.85 95.71 94.10 91.11 95.06
TABLE V: Impact of augmenting the training dataset on FR verification accuracies (authentic dataset: WF; synthetic dataset: DC). For different subsets of the authentic (WF) dataset combined with the synthetic dataset (DC), we reported the results under three settings: no augmentation, both synthetic and authentic samples are randomly augmented (RAug-All) and only synthetic samples are augmented (RAug-Synt). For each group of experimental settings, the best average accuracy is in bold. Notably, augmenting only the synthetic samples led to higher verification accuracies in most of the settings, in comparison to cases where both authentic and synthetic samples are augmented.
Train Data Method LFW ()($\uparrow$)( ↑ ) AgeDB-30 ()($\uparrow$)( ↑ ) CFP-FP ()($\uparrow$)( ↑ ) CA-LFW ()($\uparrow$)( ↑ ) CP-LFW ()($\uparrow$)( ↑ ) Avg ()($\uparrow$)( ↑ )
GC (10K) $\cup$ M2-S (1K) - 96.47 85.02 76.67 85.80 74.33 83.66
GC (10K) $\cup$ M2-S (1K) RAug-Synt 96.27 82.15 77.64 85.22 75.70 83.40
GC (10K) $\cup$ M2-S (2K) - 98.00 86.87 81.26 88.05 78.72 86.58
GC (10K) $\cup$ M2-S (2K) RAug-Synt 96.53 83.05 78.74 86.10 77.73 84.43
GC (10K) $\cup$ M2-S (4K) - 99.08 93.03 89.09 93.35 86.05 92.12
GC (10K) $\cup$ M2-S (4K) RAug-Synt 97.98 86.60 83.00 89.35 81.72 87.73
GC (10K) $\cup$ M2-S (6K) - 99.45 94.37 91.66 94.05 87.98 93.50
GC (10K) $\cup$ M2-S (6K) RAug-Synt 98.63 88.70 84.80 91.03 85.40 89.71
GC (10K) $\cup$ M2-S (8K) - 99.52 95.35 93.61 94.75 88.97 94.44
GC (10K) $\cup$ M2-S (8K) RAug-Synt 98.95 91.25 86.46 92.38 87.30 91.27
GC (10K) $\cup$ M2-S (10K) - 99.63 95.83 94.64 95.07 89.82 95.00
GC (10K) $\cup$ M2-S (10K) RAug-Synt 99.33 87.77 88.00 91.28 87.70 90.82
TABLE VI: Impact of augmenting the training dataset on FR verification accuracies (authentic dataset: M2-S; synthetic dataset: GC). Verification accuracy comparison between baselines trained with combined authentic and synthetic identities (-) and the same with data augmentation applied only on synthetic samples (RAug-Synt). For each group of results, best average accuracies are highlighted in bold. The authentic identities are sampled from M2-S, the synthetic ones from GC.
Train Data Method LFW ()($\uparrow$)( ↑ ) AgeDB-30 ()($\uparrow$)( ↑ ) CFP-FP ()($\uparrow$)( ↑ ) CA-LFW ()($\uparrow$)( ↑ ) CP-LFW ()($\uparrow$)( ↑ ) Avg ()($\uparrow$)( ↑ )
DC (10K) $\cup$ M2-S (1K) - 98.63 90.92 89.10 92.18 82.78 90.72
DC (10K) $\cup$ M2-S (1K) RAug-Synt 98.85 91.53 90.73 92.40 85.80 91.86
DC (10K) $\cup$ M2-S (2K) - 98.85 91.60 89.39 92.62 83.15 91.12
DC (10K) $\cup$ M2-S (2K) RAug-Synt 98.82 92.03 90.89 92.30 85.85 91.98
DC (10K) $\cup$ M2-S (4K) - 99.02 92.57 90.01 93.00 84.48 91.82
DC (10K) $\cup$ M2-S (4K) RAug-Synt 99.13 93.05 92.03 92.90 86.82 92.79
DC (10K) $\cup$ M2-S (6K) - 99.07 93.07 90.77 93.28 84.95 92.23
DC (10K) $\cup$ M2-S (6K) RAug-Synt 99.13 93.28 92.43 93.22 87.18 93.05
DC (10K) $\cup$ M2-S (8K) - 99.22 93.45 91.66 93.40 85.53 92.65
DC (10K) $\cup$ M2-S (8K) RAug-Synt 99.27 93.70 92.74 93.35 87.20 93.25
DC (10K) $\cup$ M2-S (10K) - 99.17 93.50 91.44 93.90 85.73 92.75
DC (10K) $\cup$ M2-S (10K) RAug-Synt 99.38 93.92 93.17 93.42 87.95 93.57
TABLE VII: Impact of augmenting the training dataset on FR verification accuracies (authentic dataset: M2-S; synthetic dataset: DC). Verification accuracy comparison between baselines trained with combined authentic and synthetic identities (-) and the same with data augmentation applied only on synthetic samples (RAug-Synt). For each group of results, best average accuracies are highlighted in bold. The authentic identities are sampled from M2-S, the synthetic ones from DC.
Given the consistent enhancements observed by augmenting only the synthetic samples in our combined datasets, it became interesting to explore whether this behavior is confirmed across different sources of authentic identities. To address this, we extended our investigation to training FR models on synthetic data combined with authentic identities from M2-S. Results are reported in Table VI and VII. In this scenario, it can be observed that combining augmented GC data with authentic identities from M2-S constantly led to worse verification performances than the combined baselines, which were impacted by $+0.23\%$, $-2.49\%$, $-4.83\%$, $-4.09\%$, $-3.38\%$ and $-4.42\%$ while involving 1K, 2K, 4K, 6K, 8K and 10K authentic identities, respectively. Combining augmented DC data with the same authentic identities instead, always led to well noticeable improvements for every authentic data subset. Specifically, FR models trained with augmented DC samples achieved $1.32\%$, $1.00\%$, $1.11\%$, $0.94\%$, $0.68\%$ and $0.93\%$ improvements. The consistence of our results while training FR models on augmented DC samples further reinforces our previous considerations, pointing to end-to-end synthetic data augmentation as a promising direction for performance improvement in FR when using combined data.
V Conclusions and Future Work
To conclude, we connect the insights coming from individual analysis and discuss the results and future research.
Our initial analysis aimed to comprehend the performance differences between FR models exclusively trained on authentic data and those trained solely on synthetic data generated through SOTA generative methods. This investigation led to a better understanding of the impact of data sources variation on model performance, under an equal number of identities across datasets. Our findings align consistently with prior research [2, 27], confirming that FR models trained on the a large, full authentic dataset exhibit the highest performance, achieving an average accuracy of $94.63\%$ across the benchmarks in our experiments.
Subsequently, we investigated whether introducing a restricted number of authentic identities led to significant gains in accuracy, compared to models exclusively trained on synthetic identities. As anticipated, our study revealed a consistent improvement across all benchmarks as the number of involved authentic identities increased. FR models trained on subsets comprising both authentic and synthetic data demonstrated remarkably competitive results when compared to models exclusively trained on the entire authentic dataset. Notably, certain combinations that included DC identities even outperformed the full authentic-only baseline.
In our third analysis, our goal was to assess whether a training dataset with a fixed number of synthetic identities and a progressively increasing number of authentic ones could attain accuracy levels, at the very least, comparable to a larger dataset comprising solely authentic identities. For both authentic datasets, all settings involving combined datasets yielded superior outcomes compared to the synthetic-only baselines, showcasing improvements of approximately $30\%$ in several cases. Moreover, training FR models on synthetic data combined with subsets of authentic identities consistently led to improvements compared to models trained exclusively on authentic subsets. As another observation, models trained on combined datasets that included synthetic data from DC consistently outperformed those involving synthetic data from GC, especially in cross-pose benchmarks.
In a last exploration, we delved into the impact of data augmentation on verification accuracy concerning the combined baselines. Throughout our experiments, we applied augmentation to two distinct sets of images: exclusively on synthetic ones and on the entire combined datasets. Notably, augmenting both synthetic and authentic samples in the combined training datasets generally resulted in inferior performance. Conversely, employing the same augmentation solely on synthetic samples led to remarkable gains in accuracy, indicating this approach as promising for enhancing accuracy when training FR models with combined datasets.
Building upon this work, our next steps involve exploring the performance of a wider range of FR backbones and incorporating additional augmentation, sampling, and domain generalization techniques. We will also explore techniques to optimize the generation of synthetic data specifically designed for training more effective FR models when faced with (very) limited authentic data, in an end-to-end manner. Additionally, in the immediate future, we plan to extend this study with a sensitivity analysis to determine an optimal parameterization for this task, encompassing both domain generalization and data augmentation techniques. This research also suggests that augmenting limited (and likely less demographically inclusive) data with more diverse synthetic data can not only improve overall performance but also enhance fairness across demographic groups. Investigating this aspect further will be one of our next research objectives.